### v1

# The Rhaly Operators on Köthe Spaces

Abstract.
We introduce and study the Rhaly operator on Köthe spaces, with a primary focus on understanding its well-definedness, continuity, and compactness. We especially examine operators acting on power series spaces of both infinite and finite type. In the sequel, we provide integral representations for the Rhaly operator on the space of entire functions $H(\mathbb{C})$ and the space of holomorphic functions on the unit disc $H(\mathbb{D})$. We also investigate the topologizability and power boundedness of the Rhaly operators, which leads to findings about their mean ergodicity, uniform mean ergodicity, and Cesáro boundedness.
Key words and phrases:
Rhaly Matrix, Köthe Spaces, Compact Operators, Power Bounded Operators
1991 Mathematics Subject Classification:
46A45, 47A05, 47B37
1. Introduction
The Cesáro operator $\mathcal{C}$ acting on the Hilbert space $\ell^{2}$ of square-summable complex numbers $a=(a_{n})_{n\in\mathbb{N}_{0}}$ is given by
$(\mathcal{C}a)(n)=\frac{1}{n+1}\sum^{n}_{k=0}a_{k}$
for $n=0,1,2,\dots$. This operator was extensively investigated in [12] and Brown, Halmos, and Shields proved, among other things, that $\mathcal{C}$ is bounded with $\|C\|_{\ell^{2}}=2$ and the spectrum $\sigma(C)=\{z:|1-z|<1\}$. Representing the Cesáro operator with respect to the standard orthonormal basis $(e_{n})_{n\in\mathbb{N}_{0}}$ for $\ell^{2}$, the Cesáro matrix
$\begin{bmatrix}1&0&0&0&\cdots\\ \frac{1}{2}&\frac{1}{2}&0&0&\cdots\\ \frac{1}{3}&\frac{1}{3}&\frac{1}{3}&0&\cdots\\ \frac{1}{4}&\frac{1}{4}&\frac{1}{4}&\frac{1}{4}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}$ (1.1)
is obtained. The Cesáro operator has been a subject of significant study within the classical sequence spaces $\ell_{p}$ ($1<p<\infty$), $\ell_{\infty}$, $c_{0}$, $c$, the weighted Banach spaces $\ell_{p}(w)$, $1<p<\infty$ and the weighted Banach spaces $c_{0}(w)$, $1<p<\infty$ for some references, we refer [1, 4, 5, 9, 12, 24].
Investigating the Cesáro operator on Fréchet spaces is also a topic of interest; we refer [6, 7]. In [7], Albanese, Bonet, and Ricker studied the Cesáro operator $\mathcal{C}$ on power series spaces of finite type $\Lambda_{1}(\alpha)$ and they focused on its spectrum, which significantly differs based on the nuclearity of $\Lambda_{1}(\alpha)$ and characterized the spectrum. Furthermore, they showed that $\mathcal{C}$ is power bounded, uniformly mean ergodic whenever $\Lambda_{1}(\alpha)$ is nuclear.
The (discrete) generalized Cesáro operators $\mathcal{C}_{t}$, $t\in\left[0,1\right]$ were first studied by Rhaly in [21], [22]. The generalized Cesáro operator $\mathcal{C}_{t}$, $t\in[0,1]$, acting on the Hilbert space $\ell^{2}$ of square-summable complex numbers $a=(a_{n})_{n\in\mathbb{N}_{0}}$ is given by
$(\mathcal{C}_{t}a)(n)=\frac{1}{n+1}\sum^{n}_{k=0}t^{n-k}a_{k}$
for $n=0,1,2,\dots$. The generalized Cesáro matrix is
$\begin{bmatrix}1&0&0&0&\cdots\\ \frac{t}{2}&\frac{1}{2}&0&0&\cdots\\ \frac{t^{2}}{3}&\frac{t}{3}&\frac{1}{3}&0&\cdots\\ \frac{t^{3}}{4}&\frac{t^{2}}{4}&\frac{t}{4}&\frac{1}{4}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}$ (1.2)
with $t\in\left[0,1\right].$ It was established in [21] that the operators $\mathcal{C}_{t}$ on $\ell^{2}$ are bounded for $0\leq t\leq 1$, and are compact for $0\leq t<1$ and has spectrum $\sigma(\mathcal{C}_{t})=\{1/n\}_{n\in\mathbb{N}}\cup\{0\}$. For generalized Cesáro operators on $c$, $c_{0}$ $\ell_{p}$ (where $1<p<\infty$), we refer to [25, 26]. Albanese, Bonet and Ricker in [8, 9] examine the continuity, compactness, spectrum, point spectrum, linear dynamics, and mean ergodicity of generalized Cesáro operators $\mathcal{C}_{t}$, $t\in[0,1]$, on the space $H(\mathbb{D})$ of holomorphic functions on the open unit disc $\mathbb{D}$, on the Banach space $H^{\infty}$ of bounded analytic functions and on the weighted Banach spaces $H^{\infty}_{\nu}$ and $H^{0}_{\nu}$ with their sup-norms, on the disc algebra $A(\mathbb{D})$, on the Hardy spaces $H^{p}$, for $1\leq p\leq\infty$.
In [23], Rhaly introduced the infinite matrices
$\begin{bmatrix}a_{0}&0&0&0&\cdots\\ a_{1}&a_{1}&0&0&\cdots\\ a_{2}&a_{2}&a_{2}&0&\cdots\\ a_{3}&a_{3}&a_{3}&a_{3}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}.$ (1.3)
This matrix is referred to as a Rhaly matrix in the literature. The classical Cesáro matrix and generalized Cesáro matrix can be obtained as a special case of a Rhaly matrix. In [23], Rhaly proved that if the limit $L=\lim_{n\to\infty}(n+1)|a_{n}|$ exists, the operator $R_{a}$ associated with this matrix is a bounded operator on $\ell^{2}$. Conversely, if this limit does not exist, $R_{a}$ is not a bounded operator on $\ell^{2}$. Furthermore, it was demonstrated that if $L=0$, the operator $R_{a}$ is compact.
In this study, we introduce and analyze the properties of Rhaly operators on Köthe spaces; Rhaly operators are defined as operators whose associated matrix is a Rhaly matrix. The author’s research also includes studies on the operators defined on Köthe spaces that are associated with Toeplitz and Hankel matrices; see [14, 15, 16] for details.
This paper is structured as follows: Section 2 provides the notational framework and aggregates the fundamental lemmas and propositions that support the main results. In Section 3, our investigation addresses the condition for the well-definedness, continuity, and compactness of Rhaly operators acting between Köthe spaces in a general setting. Specifically, Theorem 3.3 demonstrates that the Rhaly operator $R_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\alpha)$ is well-defined, continuous, and compact for every $\theta\in\Lambda_{\infty}(\alpha)$. Moreover, Theorem 3.9 details the well-definedness and continuity of the Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ for every $\theta\in\Lambda_{1}(\alpha)$. We then show in Theorem 3.10 that the Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is compact if and only if $\theta$ belongs to the dual of the nuclear power series space of finite type $\Lambda_{1}(\alpha)$. The space of all entire functions $H(\mathbb{C})$ and the space of all holomorphic functions on the unit disc $H(\mathbb{D})$ with the topology of uniform convergence on compact sets are isomorphic to the power series space of infinite type $\Lambda_{\infty}(n)$ and the power series space of finite type $\Lambda_{1}(n)$, respectively. We established the integral representations for Rhaly operators defined on the spaces $H(\mathbb{C})$ and $H(\mathbb{D})$ in section 4. The topologizability and power boundedness of the Rhaly operator on $\Lambda_{1}(\alpha)$, $\Lambda_{\infty}(\alpha)$, and for some Köthe spaces $K(a_{n,k})$ were the subject of Section 5: we demonstrated that the Rhaly operator $R_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\alpha)$ is always m-topologizable and power bounded, while the conditions for the topologizability and power boundedness of the Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ have been constructed. These findings, therefore, allowed us to determine the topologizability and power boundedness of Rhaly operators on $H(\mathbb{C})$ and $H(\mathbb{D})$. Using the facts presented in Theorem 2.6 (by Albanese, Bonet, and Ricker) and Theorem 2.7 (by Kalmes and Santacreu), we then deduce the mean ergodicity, uniform mean ergodicity, and Cesàro boundedness of Rhaly operators.
2. Preliminaries
This section provides the essential background for the paper. Our terminology and notation follow [20].
A Fréchet space is a complete Hausdorff locally convex space whose topology can be described by a countable system of seminorms $(\|\cdot\|_{k})_{k\in\mathbb{N}}$. A grading on a Fréchet space $E$ is a sequence of seminorms $\{\|\cdot\|_{n}:n\in\mathbb{N}\}$ which are increasing, that is,
$\|x\|_{1}\leq\|x\|_{2}\leq\|x\|_{3}\leq\dots$
for each $x\in E$ and which defines the topology. All Fréchet spaces admit a grading. A graded Fréchet space is a Fréchet space with a choice of grading. For more information, see [17]. In this paper, unless stated otherwise, we will assume that all Fréchet spaces are graded Fréchet spaces.
A matrix $(a_{n,k})_{k,n\in\mathbb{N}}$ of non-negative numbers is called a Köthe matrix if the following conditions hold:
  • 1.
    For each $n\in\mathbb{N}$ there exists a $k\in\mathbb{N}$ with $a_{n,k}>0$.
  • 2.
    $a_{n,k}\leq a_{n,k+1}$ for each $n,k\in\mathbb{N}$.
Given a Köthe matrix $(a_{n,k})_{n,k\in\mathbb{N}}$, the space
$K(a_{n,k})=\bigg{\{}x=(x_{n})_{n\in\mathbb{N}}\;\bigg{|}\;\|x\|_{k}:=\sum^{\infty}_{n=1}|x_{n}|a_{n,k}<\infty\;\text{for all}\;k\in\mathbb{N}\bigg{\}}$
is called a Köthe space. Every Köthe space is a Fréchet space given by the semi-norms in its definition. The dual space of a Köthe space $K(a_{n,k})$ is
$(K(a_{n,k}))^{\prime}=\bigg{\{}y=(y_{n})_{n\in\mathbb{N}}\;\bigg{|}\;\sup_{n\in\mathbb{N}}|y_{n}a_{n,k}^{-1}|<+\infty\;\text{for some}\;k\in\mathbb{N}\bigg{\}}$
see Proposition 27.3 of [20] for more information.
Grothendieck-Pietsch Criteria [20][Theorem 28.15] states that a Köthe space $K(a_{n,k})$ is nuclear if and only if for every $k\in\mathbb{N}$, there exists a $l>k$ so that
$\sum^{\infty}_{n=1}{a_{n,k}\over a_{n,l}}<\infty.$
For a nuclear Köthe space, $\|x\|_{k}=\sup_{n\in\mathbb{N}}|x_{n}|a_{n,k}$, $k\in\mathbb{N}$ forms an equivalent system of seminorms to the fundamental system of seminorms $\|x\|_{k}=\sum^{\infty}_{n=1}|x_{n}|a_{n,k}$, $k\in\mathbb{N}$.
Let $\alpha=\left(\alpha_{n}\right)_{n\in\mathbb{N}}$ be a non-negative increasing sequence with $\lim_{n\rightarrow\infty}\alpha_{n}=\infty$. The power series space of finite type associated with $\alpha$ is defined as
$\Lambda_{1}\left(\alpha\right):=\left\{x=\left(x_{n}\right)_{n\in\mathbb{N}}:\;\left\|x\right\|_{k}:=\sum^{\infty}_{n=1}\left|x_{n}\right|e^{-{1\over k}\alpha_{n}}<\infty\textnormal{ for all }k\in\mathbb{N}\right\},$
and the power series space of infinite type associated with $\alpha$ is defined as
$\Lambda_{\infty}\left(\alpha\right):=\left\{x=\left(x_{n}\right)_{n\in\mathbb{N}}:\;\left\|x\right\|_{k}:=\sum^{\infty}_{n=1}\left|x_{n}\right|e^{k\alpha_{n}}<\infty\textnormal{ for all }k\in\mathbb{N}\right\}.$
Power series spaces constitute an important class of Köthe spaces. They include the spaces of holomorphic functions on $\mathbb{C}^{d}$ and on the polydisc $\mathbb{D}^{d}$, where $\mathbb{D}$ denotes the unit disk in $\mathbb{C}$ and $d\in\mathbb{N}$:
$\mathcal{H}(\mathbb{C}^{d})\cong\Lambda_{\infty}(n^{\frac{1}{d}})\quad\quad\text{and}\quad\quad\mathcal{H}(\mathbb{D}^{d})\cong\Lambda_{1}(n^{\frac{1}{d}}).$
The nuclearity of a power series space of finite type $\Lambda_{1}(\alpha)$ and of infinite type $\Lambda_{\infty}(\alpha)$ is equivalent to conditions $\lim_{n\to\infty}\frac{\ln n}{\alpha_{n}}=0$ and $\sup_{n\in\mathbb{N}}\frac{\ln n}{\alpha_{n}}<+\infty$, respectively. We want to note that the spaces $\mathcal{H}(\mathbb{C}^{d})$ and $\mathcal{H}(\mathbb{D}^{d})$ are are nuclear for every $d\in\mathbb{N}$.
Let $E$ be a Fréchet space. A linear map $T:E\to E$ is called continuous if for every $k\in\mathbb{N}$ there exists $p\in\mathbb{N}$ and $C_{k}>0$ such that
$\|Tx\|_{k}\leq C_{k}\|x\|_{p}$
for all $x\in E$. $\mathcal{L}(E)$ denotes the vector space of all continuous linear maps from $E$ to $E$. The composition of an operator $T\in\mathcal{L}(E)$ with itself n-times is denoted by $T^{n}$. A linear map $T:E\to F$ is called compact if $T(U)$ is precompact in $F$ where $U$ is a neighborhood of zero of E.
Throughout this paper, $e_{n}$ stands for the sequence with 1 in the $n^{th}$ component and zeros elsewhere. In any Köthe space, the sequence $(e_{n})_{n\in\mathbb{N}}$ forms the canonical Schauder basis.
To establish the continuity and compactness of operators defined between Köthe spaces, we will apply Lemma 2.1 of [13].
Lemma 2.1 (Lemma 2.1, [13]).
Let $K(a_{n,k})$ and $K(b_{n,k})$ be Köthe spaces.
  • a.
    $T:K(a_{n,k})\to K(b_{n,k})$ is a linear, continuous operator if and only if for each $k\in\mathbb{N}$, there exists $m\in\mathbb{N}$ such that
    $\sup_{n\in\mathbb{N}}{\frac{\|Te_{n}\|_{k}}{\|e_{n}\|_{m}}<\infty}.$
  • b.
    If $K(b_{n,k})$ is Montel, then $T:K(a_{n,k})\to K(b_{n,k})$ is a compact operator if and only if there exists m such that for all k
    $\sup_{n\in\mathbb{N}}{\frac{\|Te_{n}\|_{k}}{\|e_{n}\|_{m}}<\infty}.$
The next proposition is Proposition 2.2 of [14], which states that the continuity condition is sufficient to ensure that linear operators defined only on the basis elements are well-defined.
Proposition 2.2.
Let $K(a_{n,k})$, $K(b_{n,k})$ be Köthe spaces and $(a_{n})_{n\in\mathbb{N}}\in K(b_{n,k})$ be a sequence. Let us define a linear map $T:K(a_{n,k})\to K(b_{n,k})$ such as
$Te_{n}=a_{n}\hskip 36.135pt\text{and}\hskip 36.135ptTx=\sum^{\infty}_{n=1}x_{n}Te_{n}$
for every $x=\sum^{\infty}_{n=1}x_{n}e_{n}$ and $n\in\mathbb{N}$. If the continuity condition
$\forall k\in\mathbb{N}\quad\quad\exists m\in\mathbb{N}\quad\quad\quad\quad\sup_{n\in\mathbb{N}}\frac{\|Te_{n}\|_{k}}{\|e_{n}\|_{m}}<\infty$
holds, then $T$ is a well-defined and continuous operator.
Let $T$ be a continuous linear operator on a Fréchet space $E$. $T^{[k]}$ denotes the k-th Cesáro mean given by
$\frac{1}{k}\sum^{k}_{m=1}T^{m}$
for every $k\in\mathbb{N}$.
Definition 2.3.
Let $E$ be a Fréchet space. An operator $T\in\mathcal{L}(E)$ is called
  • topologizable if for every $p\in\mathbb{N}$ there exists $q\in\mathbb{N}$ such that for every $k\in\mathbb{N}$ there is $M_{k,p}>0$ satisfying
    $\|T^{k}(x)\|_{p}\leq M_{k,p}\|x\|_{q}$
    for every $x\in E$.
  • m-topologizable if for every $p\in\mathbb{N}$ there exist $q\in\mathbb{N}$ and $C_{p}\geq 1$ such that
    $\|T^{k}(x)\|_{p}\leq C^{k}_{p}\|x\|_{q}$
    holds for every $k\in\mathbb{N}$ and $x\in E$.
  • power bounded if for every $p\in\mathbb{N}$ there exists $q\in\mathbb{N}$ such that
    $\|T^{k}(x)\|_{p}\leq\|x\|_{q}$
    holds for every $k\in\mathbb{N}$ and $x\in E$.
  • Cesáro bounded if for every $p\in\mathbb{N}$ there exists $q\in\mathbb{N}$ such that
    $\|T^{[k]}(x)\|_{p}\leq\|x\|_{q}$
    holds for every $k\in\mathbb{N}$ and $x\in E$.
It is straightforward that power bounded operators are m-topologizable and m-topologizable operators are topologizable. In [10] Bonet proved that every bounded operator $T\in\mathcal{L}(E)$ on a locally convex space $E$ is m-topologizable [10, Proposition 7]. Therefore, every compact operator is m-topologizable.
Żelazko [27] introduced and investigated topologizable and m-topologizable operators to establish a suitable locally convex topology on $\mathcal{L}(E)$. Bonet [10] further contributed by providing examples of non-normable locally convex spaces where all operators are topologizable, or even m-topologizable. Moreover, he proved that for a Fréchet space $E$ without a continuous norm, there is at least one operator on $E$ that is not topologizable.
Lemma 2.5 of [16] demonstrates that the notions from Definition 2.3 can be applied to Köthe spaces by examining only their basis. Lemma 2.5 of [16] is formally stated as:
Lemma 2.4.
Let $K(a_{n,k})$ be a Köthe space, and let $T:K(a_{n,k})\to K(a_{n,k})$ be a linear continuous operator.
  • i.
    $T$ is topologizable if and only if for every $p\in\mathbb{N}$ there exists $q\in\mathbb{N}$ such that for every $k\in\mathbb{N}$ there is $M_{k,p}>0$ satisfying
    $\|T^{k}(e_{n})\|_{p}\leq M_{k,p}\|e_{n}\|_{q}$
    for every $n\in\mathbb{N}$.
  • ii.
    $T$ is m-topologizable if and only if for every $p\in\mathbb{N}$ there exist $q\in\mathbb{N}$ and $C_{p}\geq 1$ such that
    $\|T^{k}(e_{n})\|_{p}\leq C_{p}^{k}\|e_{n}\|_{q}$
    holds for every $k,n\in\mathbb{N}$.
  • iii.
    $T$ is power bounded if and only if for every $p\in\mathbb{N}$ there exists $q\in\mathbb{N}$ such that
    $\|T^{k}(e_{n})\|_{p}\leq\|e_{n}\|_{q}$
    holds for every $k,n\in\mathbb{N}$.
  • iv.
    $T$ is Cesáro bounded if and only if for every $p\in\mathbb{N}$ there exists $q\in\mathbb{N}$ such that
    $\|T^{[k]}(e_{n})\|_{p}\leq\|e_{n}\|_{q}$
    holds for every $k,n\in\mathbb{N}$.
Definition 2.5.
A continuous linear operator $T$ on a locally convex Hausdorff space $E$ is called mean ergodic if the limits
$Px=\lim_{n\to\infty}\frac{1}{n}\sum^{n}_{m=1}T^{m}x,\hskip 10.84006ptx\in E$
exist in $E$. If the convergence is uniform on bounded subsets of E, then T is called uniformly mean ergodic.
A locally convex Hausdorff space E is defined as (uniformly) mean ergodic if every power bounded operator acting on it is (uniformly) mean ergodic. In this context, Albanese, Bonet, and Ricker established the following result in [2, Proposition 2.13]:
Theorem 2.6.
Let $K(a_{n,k})$ be a Köthe space. Then the following assertions are equivalent.
  • (i)
    $K(a_{n,k})$ is mean ergodic.
  • (ii)
    $K(a_{n,k})$ is uniformly mean ergodic.
  • (iii)
    $K(a_{n,k})$ is Montel.
As all power series spaces are Montel, any power bounded operator defined on power series spaces is (uniformly) mean ergodic.
Kalmes and Santacreu proved in Theorem 2.5 of [18] that for operators on Montel spaces, (uniform) mean ergodicity holds if and only if the operator is Cesàro bounded and fulfills a certain orbit growth condition. Theorem 2.5 of [18] is given below:
Theorem 2.7.
Let $E$ be a Montel space and let $T\in\mathcal{L}(E)$.
  • (a)
    $T$ is mean ergodic if and only if $T$ is uniformly mean ergodic.
  • (b)
    The following are equivalent.
    • (i)
      $T$ is Cesáro bounded and $\lim_{n\to\infty}\frac{T^{n}}{n}=0,$ pointwise in $E$.
    • (ii)
      $T$ is mean ergodic on $E$.
    • (iii)
      $T$ is uniformly mean ergodic on $E$.
3. Rhaly Operators
Let $\theta=(\theta_{n})_{n\in\mathbb{N}}$ be any sequence. The Rhaly matrix defined by $\theta$ is
$\begin{bmatrix}\theta_{1}&0&0&0&\cdots\\ \theta_{2}&\theta_{2}&0&0&\cdots\\ \theta_{3}&\theta_{3}&\theta_{3}&0&\cdots\\ \theta_{4}&\theta_{4}&\theta_{4}&\theta_{4}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}.$
We aim to define an operator $R_{\theta}:K(a_{n,k})\to K(b_{n,k})$ by taking $R_{\theta}e_{n}$ as the nth column of the above matrix, that is,
$R_{\theta}e_{n}=(0,\dots,0,\theta_{n},\theta_{n+1},\cdots)=\sum^{\infty}_{j=n}\theta_{j}e_{j}$
provided that $R_{\theta}e_{n}\in K(b_{n,k})$ for every $n=1,2,\dots$. Therefore, for every $x=\sum^{\infty}_{n=1}x_{n}e_{n}\in K(a_{n,k})$, the operator $R_{\theta}$ can be written as
$R_{\theta}x=\sum^{\infty}_{n=1}x_{n}R_{\theta}e_{n}=\sum^{\infty}_{n=1}\left(\sum^{n}_{j=1}x_{j}\right)\theta_{n}e_{n}.$ (3.1)
The operator $R_{\theta}:K(a_{n,k})\to K(b_{n,k})$ cannot be considered well-defined at this stage. This is due to our inability to guarantee the convergence of the series $\sum^{\infty}_{n=1}x_{n}R_{\theta}e_{n}$ converges in $K(b_{n,k})$ for every $x\in K(a_{n,k})$. In this paper, we aim to establish the conditions under which the operator $R_{\theta}$ is well-defined and continuous, with special emphasis on the case where one of the Köthe spaces is a power series space. Subsequently, we analyze its behavior between power series spaces.
If the operator $R_{\theta}:K(a_{n,k})\to K(b_{n,k})$ is well-defined, then $R_{\theta}(e_{1})=\theta\in K(b_{n,k})$ must necessarily hold. This indicates that assuming $\theta\in K(b_{n,k})$ is a natural and meaningful choice for such operators. Moreover, the inequalities
$\|R_{\theta}e_{n}\|_{k}=\sum^{\infty}_{j=n}|\theta_{j}|b_{j,k}\leq\sum^{\infty}_{j=1}|\theta_{j}|b_{j,k}=\|\theta\|_{k}$
guarantees that $R_{\theta}e_{n}\in K(b_{n,k})$ for all $n\in\mathbb{N}$ provided that $\theta\in K(b_{n,k})$. Proposition 2.2 leads us directly to the following proposition:
Proposition 3.1.
The operator $R_{\theta}:K(a_{n,k})\to K(b_{n,k})$ is well-defined and continuous if and only if $\theta\in K(b_{n,k})$ and the continuity condition
$\forall k\in\mathbb{N}\quad\quad\exists m\in\mathbb{N}\quad\quad\quad\quad\sup_{n\in\mathbb{N}}\frac{\|R_{\theta}e_{n}\|_{k}}{\|e_{n}\|_{m}}<\infty$
holds.
Our initial focus will be to explore the conditions required for an operator defined between the Köthe spaces $K(a_{n,k})$ and $K(b_{n,k})$ to be well-defined, continuous, and compact.
Proposition 3.2.
Let $K(a_{n,k}),K(b_{n,k})$ be Köthe spaces and let $\theta\in K(b_{n,k})$. If $\inf_{n\in\mathbb{N}}a_{n,m_{0}}>0$ for some $m_{0}\in\mathbb{N}$, then the Rhaly operator $R_{\theta}:K(a_{n,k})\to K(b_{n,k})$ is well-defined, continuous. If $K(b_{n,k})$ is Montel, then $R_{\theta}$ is compact.
Proof.
Let us assume that $\theta\in K(b_{n,k})$ and that $\inf_{n\in\mathbb{N}}a_{n,m_{0}}\geq C>0$ for some $m_{0}\in\mathbb{N}$ and $C>0$ holds. This gives us that for every $k\in\mathbb{N}$ there exists a constant $D_{k}>0$ such that
$\|R_{\theta}e_{n}\|_{k}=\sum^{\infty}_{j=n}|\theta_{j}|b_{j,k}\leq\sum^{\infty}_{j=1}|\theta_{j}|b_{j,k}\leq\|\theta\|_{k}\leq\frac{1}{C}\|\theta\|_{k}a_{n,m_{0}}\leq D_{k}a_{n,m_{0}}$
for every $n\in\mathbb{N}$ and it follows that
$\sup_{n\in\mathbb{N}}\frac{\|R_{\theta}e_{n}\|_{k}}{\|e_{n}\|_{m_{0}}}<\infty.$
Hence $R_{\theta}$ is well-defined and continuous from Proposition 2.2. If $K(b_{n,k})$ is a Montel space, $R_{\theta}$ is also compact from Lemma 2.1 as $m_{0}$ does not depend on $k$. ∎
The following theorem is a direct consequence of Proposition 3.2:
Theorem 3.3.
Let $K(b_{n,k})$ be Köthe space and let $\theta\in K(b_{n,k})$. The Rhaly operator $R_{\theta}:\Lambda_{\infty}(\alpha)\to K(b_{n,k})$ is well-defined and continuous. Furthermore, if $K(b_{n,k})$ is Montel, $R_{\theta}$ is also compact. More precisely, the following results are obtained:
  • i.
    If $\theta\in\Lambda_{\infty}(\beta)$, $R_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\beta)$ is continuous and compact.
  • ii.
    If $\theta\in\Lambda_{1}(\beta)$, $R_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{1}(\beta)$ is continuous and compact.
We now seek a weaker assumption on the matrix of $K(a_{n,k})$ in Proposition 3.2. When $K(b_{n,k})$ is a power series space of infinite type $\Lambda_{\infty}(\beta)$, a kind of stability condition on the sequence $\beta$ is required: A sequence $\beta$ is said to be weakly-stable if
$\sup_{n\in\mathbb{N}}{\beta_{n+1}\over\beta_{n}}<\infty.$
Proposition 3.4.
Let $\beta=\left(\beta_{n}\right)_{n\in\mathbb{N}}$ be a weakly-stable sequence. The Rhaly operator $R_{\theta}:K(a_{n,k})\to\Lambda_{\infty}(\beta)$ is well-defined and continuous if $\theta\in\Lambda_{\infty}(\beta)$ and the following condition holds:
$\hskip 14.45377pt\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad\quad e^{-k\beta_{n}}\leq Ca_{n,m}\hskip 43.36243pt\forall n\in\mathbb{N}.$ (3.2)
Proof.
Given that $\beta$ is a weakly-stable sequence, there exists an $M\in\mathbb{N}$, $M>1$, satisfying
$\beta_{n+1}\leq M\beta_{n}$
for every $n\in\mathbb{N}$. Let us assume that $\theta\in\Lambda_{\infty}(\beta)$ and the condition (3.2) holds. Then, for every $k\in\mathbb{N}$ there exist $m\in\mathbb{N}$, $C_{1},C_{2}>0$ such that
$\begin{split}\|R_{\theta}e_{n}\|_{k}&=\sum^{\infty}_{j=n}|\theta_{j}|e^{k\beta_{j}}=\sum^{\infty}_{j=n}|\theta_{j}|e^{k\beta_{j}}e^{k\beta_{j+1}}e^{-k\beta_{j+1}}\\ &\leq e^{-k\beta_{n+1}}\sum^{\infty}_{j=n}|\theta_{j}|e^{k(M+1)\beta_{j}}\leq e^{-k\beta_{n}}\sum^{\infty}_{j=1}|\theta_{j}|e^{k(M+1)\beta_{j}}\\ &\leq C_{1}a_{n,m}\|\theta\|_{(M+1)k}\leq C_{2}\|e_{n}\|_{m}.\end{split}$
hold for every $n\in\mathbb{N}$ and it follows that
$\sup_{n\in\mathbb{N}}\frac{\|R_{\theta}e_{n}\|_{k}}{\|e_{n}\|_{m}}<\infty.$
Hence, $R_{\theta}$ is well-defined and continuous by Proposition 2.2. ∎
By modifying the condition in (3.2), we can establish a condition for the compactness of the operator $R_{\theta}:K(a_{n,k})\to\Lambda_{\infty}(\beta)$. The proof of the following proposition has been omitted because it is similar to the proof of the proposition above.
Proposition 3.5.
Let $\beta=\left(\beta_{n}\right)_{n\in\mathbb{N}}$ be a weakly-stable sequence. The Rhaly operator $R_{\theta}:K(a_{n,k})\to\Lambda_{\infty}(\beta)$ is compact if $\theta\in\Lambda_{\infty}(\beta)$ and the following condition holds:
$\hskip 14.45377pt\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad e^{-k\beta_{n}}\leq Ca_{n,m}\hskip 43.36243pt\forall n\in\mathbb{N}.$ (3.3)
Combining Proposition 3.4 and Proposition 3.5, we conclude that the operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{\infty}(\beta)$ is well-defined, continuous, and compact, provided certain conditions on $\alpha$ and $\beta$ are met.
Theorem 3.6.
Let $\beta=\left(\beta_{n}\right)_{n\in\mathbb{N}}$ be a weakly-stable sequence. Assume that there exist $A,B>0$ such that
$\alpha_{n}\leq A\beta_{n}+B$ (3.4)
for all $n\in\mathbb{N}$. For every $\theta\in\Lambda_{\infty}(\beta)$, the Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{\infty}(\beta)$ is well-defined, continous and compact.
Proof.
Let us assume that there exist $A,B>0$ satisfying $\alpha_{n}\leq A\beta_{n}+B$ for all $n\in\mathbb{N}$. Then, for all $m,k\in\mathbb{N}$ we write
$\frac{1}{mA}\alpha_{n}-\frac{B}{A}\leq\frac{1}{A}\alpha_{n}-\frac{B}{A}\leq\beta_{n}\leq k\beta_{n}$
and
$-k\beta_{n}\leq\frac{B}{A}-\frac{1}{mA}\alpha_{n}$
for all $n\in\mathbb{N}$. Then for all $\tilde{m}\in\mathbb{N}$ satisfying $\tilde{m}>mA$ and for all $k\in\mathbb{N}$, there exists a $C>0$ such that
$e^{-k\beta_{n}}\leq Ce^{-\frac{1}{\tilde{m}}\alpha_{n}}$
for every $n\in\mathbb{N}$. This says that the conditions in (3.2) and (3.3) are satisfies. From Proposition 3.4 and Proposition 3.5, $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{\infty}(\beta)$ is well-defined, continuous and compact. ∎
We now focus on the case where the range space is a power series space of finite type.
Proposition 3.7.
Let $K(a_{n,k})$ be a Köthe space, $\Lambda_{1}(\beta)$ be a power series space of finite type. Suppose that $\theta\in\Lambda_{1}(\beta)$ and the condition
$\hskip 14.45377pt\forall k\in\mathbb{N}\quad\exists m\in\mathbb{N},C>0\quad\quad\quad e^{-\frac{1}{k}\beta_{n}}\leq Ca_{n,m}\hskip 36.135pt\forall n\in\mathbb{N}$ (3.5)
holds. Then the Rhaly operator $R_{\theta}:K(a_{n,k})\to\Lambda_{1}(\beta)$ is well-defined and continuous.
Proof.
Let us assume that $\theta\in\Lambda_{1}(\beta)$ and the condition (3.5) is satisfied. These give us that for every $k\in\mathbb{N}$, there exist $m\in\mathbb{N}$ and $C_{1},C_{2}>0$ such that
$\begin{split}\|R_{\theta}e_{n}\|_{k}&=\sum^{\infty}_{j=n}|\theta_{j}|e^{-\frac{1}{k}\beta_{j}}=\sum^{\infty}_{j=n}|\theta_{j}|e^{-\frac{1}{2k}\beta_{j}}e^{-\frac{1}{2k}\beta_{j}}\leq e^{-\frac{1}{2k}\beta_{n}}\sum^{\infty}_{j=n}|\theta_{j}|e^{-\frac{1}{2k}\beta_{j}}\\ &\leq C_{1}a_{n,m}\sum^{\infty}_{j=n}|\theta_{j}|e^{-\frac{1}{2k}\beta_{j}}=C_{1}\|\theta\|_{2k}a_{n,m}\leq C_{2}a_{n,m}\end{split}$
for every $n\in\mathbb{N}$. Therefore, for every $k\in\mathbb{N}$ there exists $m\in\mathbb{N}$ such that
$\sup_{n\in\mathbb{N}}\frac{\|R_{\theta}e_{n}\|_{k}}{\|e_{n}\|_{m}}<\infty,$
that is, $R:K(a_{n,k})\to\Lambda_{1}(\beta)$ is well defined and continuous by Proposition 2.2. ∎
We can establish a condition for the compactness of the operator $R_{\theta}:K(a_{n,k})\to\Lambda_{1}(\beta)$ by altering the condition in (3.5). The proof of the subsequent proposition has been omitted since it’s analogous to the proof of the earlier proposition.
Proposition 3.8.
Let $K(a_{n,k})$ be a Köthe space, $\Lambda_{1}(\beta)$ be a power series space of finite type. Suppose that $\theta\in\Lambda_{1}(\beta)$ and the condition
$\hskip 7.22743pt\exists m\in\mathbb{N}\quad\forall k\in\mathbb{N}\quad\exists C>0\quad\quad\quad e^{-\frac{1}{k}\beta_{n}}\leq Ca_{n,m}\hskip 36.135pt\forall n\in\mathbb{N}$ (3.6)
holds. Then the Rhaly operator $R_{\theta}:K(a_{n,k})\to\Lambda_{1}(\beta)$ is compact.
By Proposition 3.7, it follows that the operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\beta)$ is well-defined and continuous, given that specific conditions on $\alpha$ and $\beta$ hold.
Theorem 3.9.
Suppose that there exist $A,B>0$ such that
$\alpha_{n}\leq A\beta_{n}+B$ (3.7)
for all $n\in\mathbb{N}$. For every $\theta\in\Lambda_{1}(\beta)$, the Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\beta)$ is well-defined and continuous.
Proof.
Let us assume that there exist $A,B>0$ satisfying $\alpha_{n}\leq A\beta_{n}+B$ for all $n\in\mathbb{N}$. It follows that
$\frac{1}{kA}\alpha_{n}\leq\frac{1}{k}\beta_{n}+\frac{B}{A}\hskip 18.06749pt\text{and}\hskip 18.06749pt-\frac{1}{k}\beta_{n}\leq\frac{B}{A}-\frac{1}{kA}\alpha_{n}$
for every $k,n\in\mathbb{N}$. Then for all $k,m\in\mathbb{N}$ satisfying $m>kA$, there exists a $C>0$ such that
$e^{-\frac{1}{k}\beta_{n}}\leq Ce^{-\frac{1}{m}\alpha_{n}}$
for every $n\in\mathbb{N}$. This says that the condition in (3.5) is satisfied. $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\beta)$ is well-defined, continuous by Proposition 3.8. ∎
Theorem 3.10.
Let $\Lambda_{1}(\alpha)$ be a nuclear power series space of finite type and $\theta\in\Lambda_{1}(\alpha)$. The Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is compact if and only if $\theta\in(\Lambda_{1}(\alpha))^{\prime}$.
Proof.
Let us assume that the Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is compact, that is, there exists a $m_{0}\in\mathbb{N}$ so that for every $p\in\mathbb{N}$, there is a $D_{p}>0$ satisfying
$\sup_{n\in\mathbb{N}}\frac{\|R_{\theta}e_{n}\|_{p}}{\|e_{n}\|_{m_{0}}}\leq D_{p}$
and this gives us that
$|\theta_{n}|e^{\left(\frac{1}{m_{0}}-\frac{1}{p}\right)\alpha_{n}}\leq D_{p}$
for every $n\in\mathbb{N}$. We choose $p$ as $2m_{0}$. Then there exists a $D_{m_{0}}$ satisfying
$|\theta_{n}|\leq D_{m_{0}}e^{-\frac{1}{2m_{0}}\alpha_{n}}$
for every $n\in\mathbb{N}$. This says that $\theta\in(\Lambda_{1}(\alpha))^{\prime}$. For the converse, let us assume that $\theta\in(\Lambda_{1}(\alpha))^{\prime}$. There exist $r_{0}\in\mathbb{N}$ and $D_{r_{0}}>0$ so that
$|\theta_{n}|\leq D_{r_{0}}e^{-\frac{1}{r_{0}}\alpha_{n}}$
for every $n\in\mathbb{N}$. Then for every $p,n\in\mathbb{N}$, we have
$\begin{split}\frac{\|R_{\theta}e_{n}\|_{p}}{\|e_{n}\|_{r_{0}}}&=e^{\frac{1}{r_{0}}\alpha_{n}}\sup_{j\geq n}|\theta_{j}|e^{-\frac{1}{p}\alpha_{j}}\leq\sup_{j\geq n}|\theta_{j}|e^{(\frac{1}{r_{0}}-\frac{1}{p})\alpha_{j}}\\ &\leq D_{r_{0}}\sup_{j\geq n}e^{-\frac{1}{r_{0}}\alpha_{j}}e^{(\frac{1}{r_{0}}-\frac{1}{p})\alpha_{j}}\leq D_{r_{0}}.\end{split}$
This says that $R_{\theta}$ is compact. ∎
As shown by Albanese, Bonet, and Ricker in Proposition 2.5 of [7], the Cesáro operator on $\Lambda_{1}(\alpha)$ is not compact. This property is consistent with the sequence $\left(\frac{1}{n+1}\right)_{n\in\mathbb{N}_{0}}$ not being an element of the dual of $\Lambda_{1}(\alpha)$.
4. Integral Representation of Rhaly Operators
This section presents the integral representations of the Rhaly operator on the space of all entire functions $H(\mathbb{C})$ and the space of all holomorphic functions on the unit disc $H(\mathbb{D})$.
For every $g\in H(\mathbb{C})$, one can define an operator $R_{g}:H(\mathbb{C})\to H(\mathbb{C})$ such that for every $f\in H(\mathbb{C})$ and $z\in\mathbb{C}$, put
[…]
where $r_{0}$ is a constant with $0<r_{0}<1$. Since $f$ and $g$ are entire, the function $w\to\frac{f(w)}{w(1-w)}g\left(\frac{z}{w}\right)$ is holomorphic away from the origin and $w=1$, and the integration path avoids singularities, the operator is well-defined.
Theorem 4.1.
The following statements are equivalent:
$R_{\theta}:H(\mathbb{C})\to H(\mathbb{C})$ is a continuous linear operator whose associated matrix is
$\begin{bmatrix}\theta_{0}&0&0&0&\cdots\\ \theta_{1}&\theta_{1}&0&0&\cdots\\ \theta_{2}&\theta_{2}&\theta_{2}&0&\cdots\\ \theta_{3}&\theta_{3}&\theta_{3}&\theta_{3}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}.$
There exists an entire function g such that $R_{\theta}=R_{g}$. Then for every $n\in\mathbb{N}_{0}$
[…]
Proof.
Let $f(z)=\sum^{\infty}_{n=0}a_{n}z^{n}$ and $g(z)=\sum^{\infty}_{n=0}b_{n}z^{n}$ be two entire functions. We want to show that
[…]
for every $z\in\mathbb{C}$. Let us define $h(z)=\sum^{\infty}_{n=0}\left(\sum^{n}_{k=0}a_{k}\right)z^{n}$ for every $z\in\mathbb{D}$ and say $c_{n}=\sum^{n}_{k=0}a_{k}$ for every $n\in\mathbb{N}$. Then we can write that $h(z)=\sum^{\infty}_{n=0}c_{n}z^{n}$. By the following of equalities
$\begin{split}h(z)&=\sum^{\infty}_{n=0}\left(\sum^{n}_{k=0}a_{k}\right)z^{n}=\sum^{\infty}_{k=0}a_{k}\left(\sum^{\infty}_{j=k}z^{j}\right)\\ &=\sum^{\infty}_{k=0}a_{k}z^{k}\left(\sum^{\infty}_{j=0}z^{j}\right)=\frac{1}{1-z}\sum^{\infty}_{k=0}a_{k}z^{k}=\frac{1}{1-z}f(z),\end{split}$
we establish that $h(z)=\frac{f(z)}{1-z}$ for every $z\in\mathbb{D}$. Let us choose a constant $r_{0}\in(0,1)$. We obtain that
[…]
for every $z\in\mathbb{C}$. This gives us that
$(R_{g}z^{n})(\xi)=\sum^{\infty}_{m=n}b_{m}\xi^{m}$
for every $n\in\mathbb{N}_{0}$. Hence, the associated matrix of $R_{g}$ is a Rhaly matrix and
[…]
for every $n\in\mathbb{N}_{0}$. This confirms the identity $R_{\theta}=R_{g}$. On the other hand, if $R_{\theta}$, which is associated with the Rhaly matrix, is a well-defined and continuous operator, the function $g(z)=\sum^{\infty}_{n=0}\theta_{n}z^{n}$, formed by the coefficients of the sequence $\theta$, is an entire function since
$(R_{\theta}1)(z)=\sum^{\infty}_{n=0}\theta_{n}z^{n}=g(z)\in H(\mathbb{C}).$
The above calculation likewise makes it evident that $R_{\theta}=R_{g}$. ∎
As the space of all entire functions $H(\mathbb{C})$ with the topology of uniform convergence on compact subsets is isomorphic to $\Lambda_{\infty}(n)$, the following result directly stems from Theorem 3:
Corollary 4.2.
The operator $R_{g}:H(\mathbb{C})\to H(\mathbb{C})$ is compact for every $g\in H(\mathbb{C})$.
The operator $R_{g}$, which we defined on $H(\mathbb{C})$ via an integral representation, can also be defined on $H(\mathbb{D})$ as follows: For every $g\in H(\mathbb{D})$, one can define an operator $R_{g}:H(\mathbb{D})\to H(\mathbb{D})$ such that for every $f\in H(\mathbb{D})$ and $z\in\mathbb{D}$, $|z|=r_{1}$ for some $r_{1}\in(0,1)$, put
[…]
where $r_{0}$ is a constant satisfying $0<r_{0}<r_{1}<1$. Since $f,g\in H(\mathbb{D})$, the function $w\to\frac{f(w)}{w(1-w)}g\left(\frac{z}{w}\right)$ is holomorphic away from the origin, $w=1$, and the integration path avoids singularities, the operator is well-defined.
Theorem 4.3.
The following statements are equivalent:
$R_{\theta}:H(\mathbb{D})\to H(\mathbb{D})$ is a continuous linear operator whose associated matrix is
$\begin{bmatrix}\theta_{1}&0&0&0&\cdots\\ \theta_{2}&\theta_{2}&0&0&\cdots\\ \theta_{3}&\theta_{3}&\theta_{3}&0&\cdots\\ \theta_{4}&\theta_{4}&\theta_{4}&\theta_{4}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\ddots\end{bmatrix}.$
There exists a function $g\in H(\mathbb{D})$ such that $R_{\theta}=R_{g}$. Hence
[…]
We omit the proof, as it proceeds along the same lines as that of Theorem 4.1.
Given that the space of all holomorphic functions on the unit disc $H(\mathbb{D})$ with the topology of uniform convergence on compact subsets is isomorphic to $\Lambda_{1}(n)$, we can present the following result as a direct consequence of Theorem 3.10:
Corollary 4.4.
The operator $R_{g}:H(\mathbb{D})\to H(\mathbb{D})$ is compact if and only if $g$ belongs to the strong dual of $H(\mathbb{D})$.
5. Topologizability and Power Boundedness of Rhaly Operators
The topologizability and power boundedness of the Rhaly operator will be established for $\Lambda_{1}(\alpha)$, $\Lambda_{\infty}(\alpha)$, and for some Köthe spaces $K(a_{n,k})$. Consequently, these results will enable us to determine the properties of Rhaly operators on $H(\mathbb{C})$ and $H(\mathbb{D})$.
In Proposition 3.2, we showed that the operator $R_{\theta}:K(a_{n,k})\to K(b_{n,k})$ is compact if $\inf_{n\in\mathbb{N}}a_{n,m_{0}}>0$ for some $m_{0}\in\mathbb{N}$ and $K(a_{n,k})$ is Montel. As all compact operators are m-topologizable, the operator $R_{\theta}:K(a_{n,k})\to K(a_{n,k})$ is m-topologizable. In the following proposition, we will show that the operator $R_{\theta}:K(a_{n,k})\to K(a_{n,k})$ is m-topologizable without assuming that $K(a_{n,k})$ is Montel.
Proposition 5.1.
Let $K(a_{n,k})$ be a Köthe space and let $\theta\in K(a_{n,k})$. Assume that $\inf_{n\in\mathbb{N}}a_{n,m_{0}}>0$ for some $m_{0}\in\mathbb{N}$. The Rhaly operator $R_{\theta}:K(a_{n,k})\to K(a_{n,k})$ is m-topologizable.
Proof.
Let us assume that $\theta\in K(a_{n,k})$ and $A=\inf_{n\in\mathbb{N}}a_{n,m_{0}}>0$ holds for some $m_{0}\in\mathbb{N}$. For every $p\in\mathbb{N}$, we define $D_{p}:=\frac{\|\theta\|_{p}}{A}\geq 0$ and we obtain
$\|R_{\theta}e_{n}\|_{p}=\sum^{\infty}_{j=n}|\theta_{j}|a_{j,p}\leq\sum^{\infty}_{j=1}|\theta_{j}|a_{j,p}\leq\|\theta\|_{p}\leq\frac{1}{A}\|\theta\|_{p}a_{n,m_{0}}=D_{p}\|e_{n}\|_{m_{0}}$
for every $n\in\mathbb{N}$. Then it follows that
$\|R_{\theta}x\|_{p}\leq\sum^{\infty}_{n=1}|x_{n}|\|R_{\theta}e_{n}\|_{p}\leq D_{p}\sum^{\infty}_{n=1}|x_{n}|\|e_{n}\|_{m_{0}}\leq D_{p}\|x\|_{m_{0}}$
for every $x\in K(a_{n,k})$. Then, we obtain
$\|R^{2}_{\theta}x\|_{p}\leq D_{p}\|R_{\theta}x\|_{m_{0}}\leq D_{p}D_{m_{0}}\|x\|_{m_{0}}$
for every $x\in\mathbb{N}$ and
$\|R^{k}_{\theta}e_{n}\|_{p}\leq D_{p}(D_{m_{0}})^{k-1}\|e_{n}\|_{m_{0}}\leq(\max\{D_{p},D_{m_{0}}\})^{k}\|e_{n}\|_{m_{0}}$ (5.1)
for every $k,n\in\mathbb{N}$. Lemma 2.4 says that $R_{\theta}:K(a_{n,k})\to K(a_{n,k})$ is m-topologizable. ∎
Based directly on Proposition 5.1, we obtain the following result:
Corollary 5.2.
The Rhaly operator $R_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\alpha)$ is m-topologizable for every $\theta\in\Lambda_{\infty}(\alpha)$.
Similar to Proposition 5.1, a necessary condition for operator $R_{\theta}:K(a_{n,k})\to K(a_{n,k})$ to be power bounded can be given as follows:
Proposition 5.3.
Let $K(a_{n,k})$ be a Köthe space and let $\theta\in K(a_{n,k})$. Assume that $A=\inf_{n\in\mathbb{N}}a_{n,m_{0}}>0$ for some $m_{0}\in\mathbb{N}$ and $\sup_{p\in\mathbb{N}}\|\theta\|_{p}\leq A$. The Rhaly operator $R_{\theta}:K(a_{n,k})\to K(a_{n,k})$ is power bounded.
Proof.
Let $A=\inf_{n\in\mathbb{N}}a_{n,m_{0}}>0$ for some $m_{0}\in\mathbb{N}$ and $\theta\in K(a_{n,k})$ with $\sup_{p\in\mathbb{N}}\|\theta\|_{p}\leq A$. Then for every $p\in\mathbb{N}$, the numbers $D_{p}=\frac{\|\theta\|_{p}}{A}$ are less then or equal to 1. Inequality (5.1) says that $R_{\theta}$ is power bounded. ∎
Results presented in Theorem 2.6 and Theorem 2.7 lead to the following corollary:
Corollary 5.4.
Let $K(a_{n,k})$ be a Montel Köthe space and let $\theta\in K(a_{n,k})$. Assume that $A=\inf_{n\in\mathbb{N}}a_{n,m_{0}}>0$ for some $m_{0}\in\mathbb{N}$ and $\sup_{p\in\mathbb{N}}\|\theta\|_{p}\leq A$. The Rhaly operator $R_{\theta}:K(a_{n,k})\to K(a_{n,k})$ is
  • i.
    mean ergodic,
  • ii.
    uniformly mean ergodic,
  • iii.
    Cesáro bounded and $\lim_{n\to\infty}\frac{R^{n}_{\theta}}{n}=0,$ pointwise in $K(a_{n,k})$.
Theorem 5.5.
The Rhaly operator $R_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\alpha)$ is power bounded for every $\theta\in\Lambda_{\infty}(\alpha)$.
Proof.
Let $\theta\in\Lambda_{\infty}(\alpha)$. Then for every $p\in\mathbb{N}$, there exists a $q_{p}\in\mathbb{N}$ satisfying $\|\theta\|_{p}\leq e^{q_{p}\alpha_{1}}.$ We proceed by following the steps outlined in the proof of Proposition 5.1. In this case, $m_{0}$ is choosen as $q_{1}$ and $A=\inf_{n\in\mathbb{N}}a_{n,q_{1}}=e^{q_{1}\alpha_{1}}>0$. We recall that $D_{p}=\frac{\|\theta\|_{p}}{A}$ for every $p\in\mathbb{N}$ and then we get $D_{1}\leq 1$. It can also be assumed that $e^{q_{1}\alpha_{1}}\geq 1$. By using (5.1), we obtain
$\begin{split}\|R^{k}_{\theta}e_{n}\|_{p}&\leq D_{p}(D_{1})^{k-1}\|e_{n}\|_{q_{1}}\leq\frac{\|\theta\|_{p}}{e^{q_{1}\alpha_{1}}}e^{q_{1}\alpha_{n}}\\ &\leq\|\theta\|_{p}e^{q_{1}\alpha_{n}}\leq e^{(q_{p}+q_{1})\alpha_{n}}=\|e_{n}\|_{q_{p}+q_{1}}\end{split}$
By Lemma 2.4, $R_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\alpha)$ is power bounded. ∎
From Theorem 2.6 and Theorem 2.7, we present the following corollary:
Corollary 5.6.
For every $\theta\in\Lambda_{\infty}(\alpha)$, the Rhaly operator $R_{\theta}:\Lambda_{\infty}(\alpha)\to\Lambda_{\infty}(\alpha)$ is
  • i.
    mean ergodic,
  • ii.
    uniformly mean ergodic,
  • iii.
    Cesáro bounded and $\lim_{n\to\infty}\frac{R_{\theta}^{\hskip 1.8063ptn}}{n}=0,$ pointwise in $\Lambda_{\infty}(\alpha)$.
Given that $H(\mathbb{C})$ and $\Lambda_{\infty}(n)$ are isomorphic, and all relevant properties are preserved under isomorphism, the following result can be stated in conjunction with the findings presented above.
Corollary 5.7.
Let $R_{g}:H(\mathbb{C})\to H(\mathbb{C})$ is a continuous linear operator whose associated matrix is a Rhaly matrix. The operator $R_{g}$ is
  • i.
    m-topologizable,
  • ii.
    power bounded,
  • iii.
    mean ergodic,
  • iv.
    uniformly mean ergodic,
  • v.
    Cesáro bounded and $\lim_{n\to\infty}\frac{R^{n}_{g}}{n}=0,$ pointwise in $\Lambda_{\infty}(\alpha)$.
Our next step is to explore the properties of the Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$. We start by computing its powers and their corresponding seminorms. Let us assume that $\theta\in\Lambda_{1}(\alpha)$ and $\Lambda_{1}(\alpha)$ is nuclear. Since for every $n\in\mathbb{N}$ and $x\in\Lambda_{1}(\alpha)$, $R_{\theta}e_{n}=\sum^{\infty}_{j=n}\theta_{j}e_{j}$ and $R_{\theta}x=\sum^{\infty}_{n=1}\left(\sum^{n}_{i=1}x_{i}\right)\theta_{n}e_{n}$, it follows that
$R^{2}_{\theta}e_{n}=R_{\theta}(R_{\theta}e_{n})=\sum^{\infty}_{j=n}\theta_{j}\left(\sum^{j}_{i=n}\theta_{i}\right)e_{j}\hskip 10.84006pt\text{and}\hskip 10.84006pt(R^{2}_{\theta}e_{n})_{m}=\theta_{m}\sum^{m}_{j=n}\theta_{j}$
for every $m\geq n$, $n\in\mathbb{N}$. This implies that
$\|R^{2}_{\theta}e_{n}\|_{p}\leq\sup_{m\geq n}|\theta_{m}|\left(\sum^{m}_{j=n}|\theta_{j}|\right)e^{-\frac{1}{p}\alpha_{m}}$
for every $p,n\in\mathbb{N}$. In general, the coefficients of the power are written as
$(R^{k}_{\theta}e_{n})_{m}=\sum_{n\leq j_{1}\leq j_{2}\leq\cdots\leq j_{k}=m}\theta_{j_{1}}\cdots\theta_{j_{k}}$
for every $m\geq n$, $k,n\in\mathbb{N}$. Hence, we obtain that
$|(R^{k}_{\theta}e_{n})_{m}|\leq\left(\sum^{m}_{j=n}|\theta_{j}|\right)^{k}\hskip 10.84006pt\text{and}\hskip 10.84006pt\|R^{k}_{\theta}e_{n}\|_{p}\leq\sup_{m\geq n}\left(\sum^{m}_{j=n}|\theta_{j}|^{k}\right)e^{-\frac{1}{p}\alpha_{m}}$
for every $m\geq n$, $n,k,p\in\mathbb{N}$. Consequently, the inequalities
$\begin{split}\|R^{k}_{\theta}e_{n}\|_{p}&\leq\sup_{m\geq n}\left(\sum^{m}_{j=n}|\theta_{j}|^{k}\right)e^{-\frac{1}{p}\alpha_{m}}\\ &=\sup_{m\geq n}\left(\sum^{m}_{j=n}|\theta_{j}|^{k}\right)e^{-\frac{1}{3p}\alpha_{m}}e^{-\frac{1}{3p}\alpha_{m}}e^{-\frac{1}{3p}\alpha_{m}}\\ &\leq e^{-\frac{1}{3p}\alpha_{n}}\sup_{m\geq n}\left(\sum^{m}_{j=n}|\theta_{j}|e^{-\frac{1}{3pk}\alpha_{j}}\right)^{k}e^{-\frac{1}{3p}\alpha_{m}}\\ &\leq e^{-\frac{1}{3p}\alpha_{n}}\left(\sum^{\infty}_{j=n}|\theta_{j}|e^{-\frac{1}{3pk}\alpha_{j}}\right)^{k}\sup_{m\geq n}e^{-\frac{1}{3p}\alpha_{m}}\\ &\leq e^{-\frac{1}{3p}\alpha_{n}}\left(\|\theta\|_{3pk}\right)^{k}=\|e_{n}\|_{3p}\left(\|\theta\|_{3pk}\right)^{k}\end{split}$ (5.2)
are satisfied for every $n,k,p\in\mathbb{N}$. As a direct corollary of the inequality (5.2) and Lemma 2.4, we obtain the following:
Proposition 5.8.
Let $\Lambda_{1}(\alpha)$ be a nuclear space and $\theta\in\Lambda_{1}(\alpha)$. The Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is topologizable. If $\sup_{p\in\mathbb{N}}\|\theta\|_{p}<\infty$, $R_{\theta}$ is m-topologizable.
Theorem 5.9.
Let $\Lambda_{1}(\alpha)$ be a nuclear space and $\theta\in\Lambda_{1}(\alpha)$. The Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is power bounded if and only if $\sup_{p\in\mathbb{N}}\|\theta\|_{p}\leq 1$.
Proof.
Let assume that $\sup_{p\in\mathbb{N}}\|\theta\|_{p}\leq 1$. (5.2) gives us that for every $p,k,n\in\mathbb{N}$, it yields that
$\|R^{k}_{\theta}e_{n}\|_{p}\leq\|e_{n}\|_{3p}.$
Then $R_{\theta}$ is power bounded. For the converse, we assume that $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$ is power bounded. This means that for every $p\in\mathbb{N}$ there exists a $q\in\mathbb{N}$ such that
$\|R^{k}_{\theta}e_{n}\|_{p}\leq\|e_{n}\|_{q}$
holds for every $k,n\in\mathbb{N}$. By choosing $n=k=1$, we have
$\|\theta\|_{p}=\|R_{\theta}e_{1}\|_{p}\leq\|e_{1}\|_{q}=e^{-\frac{1}{q}\alpha_{1}}<1.$
This completes the proof. ∎
Building upon Theorem 2.6 and Theorem 2.7, we establish the following result:
Corollary 5.10.
Let $\Lambda_{1}(\alpha)$ be a nuclear power series space of finite type. The following properties are equivalent for the Rhaly operator $R_{\theta}:\Lambda_{1}(\alpha)\to\Lambda_{1}(\alpha)$, where $\theta\in\Lambda_{1}(\alpha)$;
  • 1.
    $R_{\theta}$ is power bounded.
  • 2.
    $R_{\theta}$ mean ergodic.
  • 3.
    $R_{\theta}$ is uniformly mean ergodic.
  • 4.
    $R_{\theta}$ is Cesáro bounded and $\lim_{n\to\infty}\frac{R_{\theta}^{\hskip 1.8063ptn}}{n}=0,$ pointwise in $\Lambda_{1}(\alpha)$.
  • 5.
    $\sup_{p\in\mathbb{N}}\|\theta\|_{p}\leq 1$
Proof.
Theorem 5.9 asserts the equivalence between conditions (1) and (5). Theorem 2.6 (by Albanese, Bonet, and Ricker) regarding the mean ergodicity of Montel Köthe spaces, establishes that (1) implies (2). Furthermore, Theorem 2.7 (by Kalmes and Santacreu) tells us that statements (2), (3), and (4) are equivalent. We now turn to proving that (4) implies (5). Let us assume that $R_{\theta}$ is Cesáro bounded. Then, for every $p\in\mathbb{N}$, there exists a $q\in\mathbb{N}$ such that for every $k,n\in\mathbb{N}$
$\|R^{[k]}_{\theta}(e_{n})\|_{p}\leq\|e_{n}\|_{q}$
holds. In particular, taking $k=n=1$, we get
$\|\theta\|_{p}=\|R_{\theta}(e_{1})\|_{p}\leq\|e_{1}\|_{q}=e^{-\frac{1}{q}\alpha_{1}}<1$
which shows that (5) is satisfied. This completes the proof. ∎
Since $H(\mathbb{D})$ and $\Lambda_{1}(n)$ are isomorphic, and all properties of interest remain invariant under isomorphism, we can state the following result in connection with the results given above.
Corollary 5.11.
Let $R_{g}:H(\mathbb{D})\to H(\mathbb{D})$ be a continuous linear operator whose associated matrix is a Rhaly matrix. The operator $R_{g}$ is
  • a.
    topologizable
  • b.
    m-topologizable if $\sup_{p\in\mathbb{N}}\|g\|_{p}<+\infty$.
The following statements are equivalent for the operator $R_{\theta}:H(\mathbb{D})\to H(\mathbb{D})$:
  • i.
    power bounded,
  • ii.
    mean ergodic,
  • iii.
    uniformly mean ergodic,
  • iv.
    Cesáro bounded and $\lim_{n\to\infty}\frac{R^{n}_{g}}{n}=0,$ pointwise in $\Lambda_{\infty}(\alpha)$.
  • v.
    $\sup_{p\in\mathbb{N}}\|g\|_{p}\leq 1$